OpenStudy (anonymous):

y(y + 4) - y = 6 is a quadratic equation. True False

OpenStudy (anonymous):

A quadratic equation is an equation of the form: ax^2 + bx + c = 0

OpenStudy (anonymous):

i sill dont get it can you explian

OpenStudy (anonymous):

how do i know if this problem is true or false

OpenStudy (anonymous):

y( y + 4) - y = 6 y^2 + 4y - y = 6 y^2 + 3y = 6 y^2 + 3y - 6 = 0 <---Yes (:

OpenStudy (anonymous):

thanks